The following table gives you genotype frequencies at two blood loci in humans

A) Estimate the allele frequencies at each locus

B) Test for Hardy-Weinberg equilibrium at each locus

C) Estimate D and D'

D) Test for significant gametic disequilibrium

MM MN NN
SS 91 32 5
Ss 147 78 17
ss 82 75 7
Answer #1

(A) MM7 (91+147+82)/3 = 106.67 (apbook) MN =(32+28+75) /+61-67(abbrox) NN (5+17++)/3 = 25 SS +(91+32+5)/3 = 42.67(abbrox) So=